OpenStudy (anonymous):

whats the volume of y=x^(2/3) , x=1, y=0; about the y-axis?

OpenStudy (anonymous):

check ur question again plz

OpenStudy (anonymous):

whats the volume of y=x^(2/3) , x=1, y=0; rotated about the y-axis?

OpenStudy (anonymous):

i got 3pi/4, but i'm not sure

OpenStudy (anonymous):

hmm i got pi/4 the integral part i had i think is wrong i got from 0 to 1

OpenStudy (anonymous):

actually in this case, xrun from 0 to 1, y run from 0 to x^(2/3), so i think we can calculate the area, then multiply by 2pi to find volume

OpenStudy (anonymous):

using double integral

OpenStudy (anonymous):

hmm i think im getting it

OpenStudy (anonymous):

hey dude, why u mean by x=1, y=0??? Are these a point or a vector???

OpenStudy (anonymous):

hey dude, what u mean by x=1, y=0??? Are these a point or a vector???

OpenStudy (anonymous):

i dont know for sure. i think points

OpenStudy (anonymous):

actually those numbers are used to find the \[?\le x \le ?\]

OpenStudy (anonymous):

I rewrote as functions of x and used the disk method. \[\pi \int\limits_{0}^{1} (1)^2-(y^(2/3))^2 dy\] \[=\pi \int\limits_{0}^{1} 1-y^3 dy\] \[= 2\pi/3\] But, I wouldn't bet my life on it.

OpenStudy (anonymous):

Actually, I should have said the washer method.

OpenStudy (anonymous):

I made a mistake evaluating the integral. The answer should be \[3\pi/4\]